# Multi-finger Manipulation via Trajectory Optimization with Differentiable Rolling and Geometric Constraints

### v1

@@ Abstract
Parameterizing finger rolling and finger-object contacts in a differentiable manner is important for formulating dexterous manipulation as a trajectory optimization problem. In contrast to previous methods which often assume simplified geometries of the robot and object or do not explicitly model finger rolling, we propose a method to further extend the capabilities of dexterous manipulation by accounting for non-trivial geometries of both the robot and the object. By integrating the object’s Signed Distance Field (SDF) with a sampling method, our method estimates contact and rolling-related variables and includes those in a trajectory optimization framework. This formulation naturally allows for the emergence of finger-rolling behaviors, enabling the robot to locally adjust the contact points. Our method is tested in a peg alignment task and a screwdriver turning task, where it outperforms the baselines in terms of achieving desired object configurations and avoiding dropping the object. We also successfully apply our method to a real-world screwdriver turning task, demonstrating its robustness to the sim2real gap.
@@ I INTRODUCTION
Multi-finger dexterous manipulation can be used to accomplish a wide range of useful dexterous manipulation tasks, especially those involving tool use, such as turning a screwdriver or orienting a part for insertion. Many methods have approached these kinds of tasks via reinforcement learning (RL) [1, 2, 3]. But while these methods are capable of producing the desired behavior given extensive training data on the specific task, they provide no guarantees on constraint satisfaction (e.g. that the fingers stay in contact). They are also ill-suited for dynamic task assignment, as a large amount of data must be collected for each new task.
@@
Figure 1: Rolling the finger can extend the flexibility of the robot and is necessary for many dexterous manipulation tasks such as screwdriver turning and peg alignment. Parameterizing the geometry of both the robot fingers and the manipulated object is important for formulating the finger-object contacts and rolling behaviors. We propose a sampling method to approximate the geometry of the robots. Integrating it with the SDF of the object, contact-related variables such as distance between meshes, contact points, and contact normals can be estimated in a differentiable manner.
There are two types of trajectory optimization that could be applied to this problem: sampling-based [7, 8] and gradient-based [9, 10]. While the sampling-based methods are more flexible, as they do not require differentiable dynamics or cost functions, they do not perform well under stringent constraints (e.g. ensuring that fingers remain in contact). Thus we choose a gradient-based trajectory optimization method, Constrained Stein Variational Trajectory Optimization (CSVTO) [9]. We choose this method because it decouples the gradient of the cost from the gradient of the constraints. This allows for guarantees on constraint satisfaction, even for low-dimensional constraint manifolds, such as those induced by contact.
@@
To overcome these challenges, we present an optimization formulation for dexterous manipulation that consists of three primary contributions:
@@
    We propose a method based on signed distance fields (SDFs) that allows us to compute gradients for contacts between arbitrary shapes.
@@
    We formulate differentiable constraints for 3D finger-rolling, which allow significant freedom in the contact interaction, enabling the method to accomplish large object orientation changes without breaking contact.
  3. 3.
    We formulate the problem quasi-statically by explicitly reasoning about force balance in a manner similar to collocation, which keeps the optimization problem tractable while ensuring the object is not dropped.
We incorporate our optimization method into an MPC framework and test our method on two multi-finger dexterous manipulation tasks: aligning a peg while maintaining contact with an external object and turning a screwdriver. Our simulated and real-world results suggest that formulating differentiable contact and finger-rolling constraints can significantly extend the capabilities of dexterous manipulation.
@@ II-A Planning for Dexterous Manipulation
There is a long history of developing planning methods for dexterous manipulation [11, 12, 13]. Many of these methods explicitly reason about changes in contact state. For instance, Xu et al. [12] propose sample-based motion planning in a hybrid configuration space for finger gait planning. Recently, the CMGMP algorithm was proposed by Cheng et al. [14] which automatically generates contact modes in a sample-based motion planning framework. Optimization-based approaches have also been used for dexterous manipulation [15, 16, 17, 18]. However, these methods do not explicitly consider the non-trivial geometries of the robot fingers and fail to account for finger contact and rolling behaviors in a differentiable manner.
@@ II-B Finger Rolling
There have been many works exploiting rolling contact for dexterous manipulation [19, 20, 21, 22, 23, 24]. Early methods were limited to the case of a spherical object rolling on a planar surface [20, 21], or with semi-spherical fingertips [22]. Bai and Liu [23] proposed a method that allows for rolling contacts with more general polyhedral geometries but is limited to non-prehensile tasks. Recent work by Tang et al. [25] proposed a method for turning a screwdriver with a dexterous hand, that also exploits rolling contacts, however, this approach is highly specialized to the task. In contrast, our proposed method allows rolling contact with arbitrary finger geometries and object geometries consisting of the composition of primitives. We also demonstrate both a screwdriver-turning task and a peg alignment task.
@@ III PROBLEM STATEMENT
In this work, we focus on the task of dexterous manipulation with a multi-fingered hand. We consider a discrete-time trajectory optimization problem with $\mathbf{s}_{t}$ as the state variable, and $\mathbf{u}_{t}$ as the action. The state variable $\mathbf{s}_{t}$ includes the robot configuration $\mathbf{q}_{t}$, and object configuration: $\mathbf{o}_{t}:=\{\bm{\theta}_{t},\mathbf{x}_{t}\}$, which consists of object position $\mathbf{x}_{t}$ and object orientation $\bm{\theta}_{t}$. We assume the existence of a low-level controller (i.e. a PD controller) responsible for executing position commands and define the action space as the delta actions, which means $\mathbf{q}_{t}+\mathbf{u}_{t}$ is the commanded next robot configuration. Our objective is to reach a desired object configuration $\mathbf{o_{g}}$. We model our problem in a trajectory optimization framework, aiming to optimize for a sequence of actions, such that the objective $J(\mathbf{\tau})$ is minimized, where $\mathbf{\tau}:=\{\mathbf{s}_{1},\mathbf{u}_{1},\cdots,\mathbf{s}_{T},\mathbf{u}_{T}\}$.
@@
However, solving a general dexterous manipulation problem without additional assumptions is extremely challenging due to the high dimensionality of the trajectory and the non-differentiable nature of contact dynamics. To make the problem tractable, we assume (1) the system is quasi-static, (2) the contact modes are prespecified and (3) the fingers are already in contact with the object at the initial step.
@@
The contact mode is defined as the type of interaction between two objects [26], e.g., sliding contact, rolling contact, and not in contact. Contact points are not created or removed within a contact mode. Specifically, in our problem, we assume a three-finger rolling contact, and the robot hand initially grasps the object with its fingertips. It also maintains this grasp throughout the manipulation process, without breaking or establishing new contacts. Despite having a fixed contact mode, our method still gives the robot the flexibility of adjusting the contacts by rolling the fingertips.
Figure 2: Rolling vs. sliding. Rigid bodies R and O are contacting at the red point $\mathbf{c}$. The contact points on R and O have the velocity $\mathbf{v}_{c,r}$ and $\mathbf{v}_{c,o}$ respectively. Fig.(a) shows pure rolling, where the contact point pair has the same velocity: $\mathbf{v}_{c,r}=\mathbf{v}_{c,o}$. Fig.(b) includes sliding, which means $\mathbf{v}_{c,r}\neq\mathbf{v}_{c,o}$. Sliding will also induce sliding friction $f_{slide}$.
In our formulation, pure rolling means the contact points share the same velocities. Sliding occurs when $\mathbf{v}_{c,r}\neq\mathbf{v}_{c,o}$ (See Fig. 2).To incorporate finger rolling into our optimization, we further assume knowledge of the geometry of both the robot fingers and the manipulated object. It is not trivial to model finger rolling since modeling of arbitrary finger and object shapes and extracting meaningful gradients based on the geometries can be very challenging.
@@ IV METHOD
In this section, we describe our trajectory optimization formulation which considers the geometries of both the fingers and the object for contact and rolling constraints. There are generally two key questions: (1) how to formulate a geometry-aware trajectory optimization problem that considers finger rolling and (2) how to process the geometries of both the robot and the object and formulate the contact and rolling constraints differentiably.
@@ IV-A Trajectory Optimization with 3D Finger Rolling
We focus on trajectory optimization with fixed contact modes. At the beginning of the task, the robot fingers are already in contact with the object. During manipulation, the robot can only utilize rolling to adjust the contact points locally but is not allowed to break and establish new contacts, such as through regrasping. We do not consider finger sliding. We also assume the system is quasi-static.
Our trajectory optimization problem can be written as:
$\min_{\begin{subarray}{c}\mathbf{s}_{1},\mathbf{s}_{2},\cdots,\mathbf{s}_{T};\\ \mathbf{u}_{1},\mathbf{u}_{1},\cdots,\mathbf{u}\end{subarray}_{T}}J_{goal}(\bm{\tau})+J_{smooth}(\bm{\tau})$ (1)
@@
$\mathbf{u}_{min}\leq\mathbf{u}_{t}\leq\mathbf{u}_{max}$
$f_{contact}(\mathbf{s}_{t})=0$
$f_{kinematics}(\mathbf{s}_{t},\mathbf{s}_{t+1})=0$
@@
$f_{friction}(\mathbf{s}_{t},\mathbf{u}_{t})\leq 0,$
where the state space is defined as $\mathbf{s}_{t}:=\{\mathbf{q}_{t},\mathbf{o}_{t}\}$. The action space is defined as $\mathbf{u}_{t}:=\{\hat{\mathbf{u}}_{t},\mathbf{f}_{1,t},\cdots,\mathbf{f}_{N,t},\mathbf{f}_{e,t}\}$, where $\hat{\mathbf{u}}_{t}$ is the delta action, $\mathbf{f}_{i,t}$ and $\mathbf{f}_{e,t}$ are the forces that the $i$th finger and the environment apply to the object respectively. The objective $J_{goal}$ incentivizes the robot to manipulate the object to the desired pose, and $J_{smooth}$ promotes a smooth trajectory. $f_{contact}$ ensures the fingers are always in contact with the object. $f_{kinematics}$, $f_{balance}$, and $f_{friction}$ are kinematics constraints, wrench balance constraints, and friction constraints, which are explained in the section below. Our primary contribution lies in the formulation of $f_{contact}$ and $f_{kinematics}$, utilizing SDF and a sampling-based method for estimating constraints and their gradients. To solve the optimization problem, we use CSVTO [9] as the solver.
@@ IV-A1 Contact Constraints
The constraints are formulated as
$\Phi_{i}(\mathbf{q}_{t},\mathbf{o}_{t})=0,$ (2)
where $\Phi_{i}$ returns the distance between the $i$th finger and the object. The distance between two shapes is defined as the minimal distance between any two points sampled from their surfaces. As robot fingers can have non-primitive geometries, estimating the constraints requires parameterizing the geometries. Additionally, the estimation must be differentiable to allow the gradients of the constraints to be used for updating the trajectory. We will further discuss the geometry processing in Sec. IV-B.
@@ IV-A2 Kinematics Constraint
Figure 3: (a) Projection of the kinematics constraint into the tangential space of the contact. Specifically, the constraint is satisfied if the blue and the green vector both lie on the red dashed line. (b) Variables used in the force balance. The robot wants to move from $\mathbf{q}_{t}$ to $\mathbf{q}_{t}+\hat{\mathbf{u}}_{t}$ but ends up at $\mathbf{q}_{t+1}$. Thus there will be an end effector force $\mathbf{f}_{i,t}$ approximately pointing in the $\mathbf{q}_{t}+\hat{\mathbf{u}}_{t}$ direction. Those forces from each finger combined with external forces will form a wrench balance.
@@
$J_{\mathbf{c}_{i,t}}(\mathbf{q}_{t})\dot{\mathbf{q}}_{t}=\bm{\omega}_{t}\times\mathbf{r}(\mathbf{c}_{i,t},\mathbf{o}_{t})+\dot{\mathbf{x}}_{t},$ (3)
@@
However, having both the kinematics constraint and the contact constraint may over-constrain the problem: both the contact constraint (1D constraint) and the kinematics constraint (3D constraint) specify the contact point location (3D variable), adding up to a 4D constraint. Thus, we project the 3D kinematics constraints into the 2D tangent space of the contacts: $R(\mathbf{n}_{i,t})(J_{\mathbf{c}_{i,t}}(\mathbf{q}_{t})\dot{\mathbf{q}}_{t}-\bm{\omega}_{t}\times\mathbf{r}(\mathbf{c}_{i,t},\mathbf{o}_{t})-\dot{\mathbf{x}}_{t}),$ where $R(\mathbf{n}_{i,t})$ is the projection matrix based on the contact normal $\mathbf{n}_{i,t}$ of the $i$th finger. The contact normal is a function of $\mathbf{s}_{t}$. The visualization of the projection is shown in Fig. 3.
Similar to contact constraints, estimating contact points $\mathbf{c}_{i,t}$, contact jacobian $J_{\mathbf{c}_{i,t}}$ and contact normal $\mathbf{n}_{i,t}$ and their gradients also requires parameterizing the geometries, which is discussed in Sec. IV-B.

### v2

@@ Abstract
Parameterizing finger rolling and finger-object contacts in a differentiable manner is important for formulating dexterous manipulation as a trajectory optimization problem. In contrast to previous methods which often assume simplified geometries of the robot and object or do not explicitly model finger rolling, we propose a method to further extend the capabilities of dexterous manipulation by accounting for non-trivial geometries of both the robot and the object. By integrating the object’s Signed Distance Field (SDF) with a sampling method, our method estimates contact and rolling-related variables in a differentiable manner and includes those in a trajectory optimization framework. This formulation naturally allows for the emergence of finger-rolling behaviors, enabling the robot to locally adjust the contact points. To evaluate our method, we introduce a benchmark featuring challenging multi-finger dexterous manipulation tasks, such as screwdriver turning and in-hand reorientation. Our method outperforms baselines in terms of achieving desired object configurations and avoiding dropping the object. We also successfully apply our method to a real-world screwdriver turning task and a cuboid alignment task, demonstrating its robustness to the sim2real gap.
@@ I INTRODUCTION
Multi-finger dexterous manipulation can be used to accomplish a wide range of useful dexterous manipulation tasks, such as turning a screwdriver or orienting an object. Many methods have approached such tasks via reinforcement learning (RL) [1, 2, 3]. While these methods are capable of producing the desired behavior given extensive task-specific training, they provide no guarantees on constraint satisfaction (e.g. maintaining finger contact). They are also ill-suited for dynamic task assignment, as a large amount of data must be collected for each new task.
@@
Figure 1: Rolling the finger can extend the flexibility of the robot and is necessary for many dexterous manipulation tasks. Parameterizing the geometry of both the robot fingers and the manipulated object is important for formulating the finger-object contacts and rolling behaviors. We propose a sampling method to approximate the geometry of the robots. Integrating it with the SDF of the object, contact-related variables such as distance between meshes, contact points, and contact normals can be estimated in a differentiable manner.
There are two types of trajectory optimization that could be applied to this problem: sampling-based [7, 8] and gradient-based [9, 10]. While the sampling-based methods are more flexible, as they do not require differentiable dynamics or cost functions, they do not perform well under stringent constraints (e.g. maintaining finger contact). Thus we choose a gradient-based method: Constrained Stein Variational Trajectory Optimization(CSVTO) [9]. This method decouples the cost and constraint gradients, ensuring constraint satisfaction even in low-dimensional manifolds like those induced by contact.
@@
To address these challenges, we make 2 key contributions:
@@
    We formulate differentiable constraints for 3D finger-rolling and finger-object contacts, which consider the non-trivial geometry of robot fingers. This differentiable formulation enables trajectory optimization for sensitive manipulation tasks and allows significant freedom in contact interaction, enabling the method to use finger rolling to accomplish large object orientation changes without breaking contact.
@@
    We propose a benchmark for multi-finger manipulation tasks, requiring the robot to adhere to stringent contact constraints to reach the goal and prevent irrecoverable failures, such as dropping the manipulated object.
We evaluate our method on our benchmark which features object reorientation tasks both with and without external environment-object contacts. Our simulated and real-world results suggest that formulating differentiable contact and finger-rolling constraints can significantly extend the capabilities of dexterous manipulation. Please see https://sites.google.com/umich.edu/multi-finger-rolling/home for more details about experiments, the appendix, and the benchmark.
@@ II-A Planning for Dexterous Manipulation
There is a long history of developing planning methods for dexterous manipulation [11, 12, 13, 14]. Many of these methods explicitly reason about changes in contact state. For instance, Xu et al. [12] propose sample-based motion planning in a hybrid configuration space for finger gait planning. Cheng et al. propose an algorithm [15], which automatically generates contact modes in a sample-based framework. Optimization-based approaches have also been used for dexterous manipulation [16, 17, 18, 19, 20, 21]. However, these methods do not explicitly consider the non-trivial geometries of the robot fingers, i.e., they approximate the robot fingertips using either a point or basic primitives such as a sphere and lack explicit differentiable modeling of finger contact or rolling behavior. Pang et al. [22] utilizes Drake [23] to compute gradients of the contact and rolling constraints. However, their implementation simplifies all geometries to basic primitives to address the discontinuity of contact-related gradients. To the best of our knowledge, this work is the first to incorporate non-trivial robot finger geometry in multi-finger dexterous manipulation tasks.
@@ II-B Finger Rolling
There have been many works exploiting rolling contact for dexterous manipulation [24, 25, 26, 27, 28, 29]. Early methods were limited to the case of a spherical object rolling on a planar surface [25, 26], or with semi-spherical fingertips [27]. Bai and Liu [28] proposed a method that allows for rolling contacts with more general polyhedral geometries but it is limited to non-prehensile tasks. Recent work by Tang et al. [30] proposed a method for turning a screwdriver with a dexterous hand, that also exploits rolling contacts, however, this approach is highly specialized to the task. In contrast, our proposed method allows rolling contact with arbitrary finger geometries and object geometries consisting of a composition of primitives. We also show our method performing five distinct tasks.
II-C Benchmarks for Multi-finger Dexterous Manipulation
Existing benchmarks for multi-finger manipulation include Chen et al.[31], which focuses on human-like bimanual manipulation; Bao et al.[32], which emphasizes articulated objects; and Cruciani et al. [33], which solely targets in-hand reorientation. However, no benchmark addresses fine multi-finger manipulation requiring precise control to prevent irrecoverable failures, such as dropping the object.
@@ III PROBLEM STATEMENT
In this work, we address dexterous manipulation with a multi-fingered hand, modeled as a discrete-time trajectory optimization problem, with the state $\mathbf{s}_{t}$, and the action $\mathbf{u}_{t}$. The state $\mathbf{s}_{t}$ includes the robot configuration $\mathbf{q}_{t}$, and object configuration: $\mathbf{o}_{t}:=\{\bm{\theta}_{t},\mathbf{x}_{t}\}$, which consists of object position $\mathbf{x}_{t}$ and object orientation $\bm{\theta}_{t}$. We assume the existence of a low-level controller (i.e. a PD controller) responsible for executing position commands and define the action space as the delta actions, which means $\mathbf{q}_{t}+\mathbf{u}_{t}$ is the commanded next robot configuration. The objective is to reach a desired object configuration $\mathbf{o_{g}}$. We model our problem in a trajectory optimization framework, aiming to optimize for a trajectory $\mathbf{\tau}:=\{\mathbf{s}_{1},\mathbf{u}_{1},\cdots,\mathbf{s}_{T},\mathbf{u}_{T}\}$, such that the objective $J(\mathbf{\tau})$ is minimized.
@@
However, solving a general dexterous manipulation problem without additional assumptions is extremely challenging due to the high dimensionality of the trajectory and the non-differentiable nature of contact dynamics. To make the problem tractable, we assume: (1) the system is quasi-static, (2) the contact modes are prespecified, and (3) the fingers are already in contact with the object initially. We further assume (4) knowledge of the geometry of robot fingers.
@@
The contact mode is defined as the type of interaction between two objects [34], e.g., sliding, rolling, and not in contact. Pure rolling occurs when the contact point velocities of the robot and object are equal ( $\mathbf{v}_{c,r}=\mathbf{v}_{c,o}$, Fig. 2(a)) Sliding occurs when $\mathbf{v}_{c,r}\neq\mathbf{v}_{c,o}$. Contact points are not created or removed within a contact mode. Specifically, in our problem, we assume a three-finger rolling contact and the robot hand initially grasps the object with its fingertips. It also maintains this grasp throughout the manipulation.
Despite having a fixed contact mode, our method still gives the robot the flexibility of adjusting the contacts by rolling its fingertips. It is not trivial to model finger rolling since modeling arbitrary finger and object shapes and extracting meaningful gradients based on the geometries can be very challenging.
Figure 2: (a) Pure Rolling. The blue(robot) and the green(object) rigid bodies contact at the red point $\mathbf{c}$. The contact points have velocities $\mathbf{v}_{c,r}$ and $\mathbf{v}_{c,o}$ respectively. Pure rolling happens when the contact point pair has the same velocity: $\mathbf{v}_{c,r}=\mathbf{v}_{c,o}$. (b) Projection of the kinematics constraint into the tangential space of the contact. Specifically, the constraint is satisfied if the blue and the green vector both lie on the red dashed line. (c) Variables used in the force balance. The robot wants to move from $\mathbf{q}_{t}$ to $\mathbf{q}_{t}+\hat{\mathbf{u}}_{t}$ but ends up at $\mathbf{q}_{t+1}$. Thus there will be an end effector force $\mathbf{f}_{i,t}$ approximately pointing in the $\mathbf{q}_{t}+\hat{\mathbf{u}}_{t}$ direction.
@@ IV METHOD
In this section, we describe our trajectory optimization formulation which considers the geometries of both the fingers and the object for contact and rolling constraints. The key questions are: (1) how to formulate a geometry-aware trajectory optimization problem that considers finger rolling and (2) how to process the geometries of both the robot and the object to formulate the contact and rolling constraints differentiably. Unlike previous methods [22, 20], our method accounts for the non-trivial geometry of robot fingers, enabling the formulation of finger rolling behavior to effectively tackle sensitive reorientation tasks, which demand precise manipulation and control.
@@ IV-A Trajectory Optimization with 3D Finger Rolling
We focus on trajectory optimization with fixed contact modes. The robot fingers start the task in contact with the object. During manipulation, the robot can only utilize rolling to adjust the contact points locally but is not allowed to break and establish new contacts, such as through regrasping. We do not consider finger sliding. We also assume the system is quasi-static. The trajectory optimization is formulated as:
$\min_{\begin{subarray}{c}\mathbf{s}_{1},\mathbf{s}_{2},\cdots,\mathbf{s}_{T};\\ \mathbf{u}_{0},\mathbf{u}_{1},\cdots,\mathbf{u}\end{subarray}_{T-1}}J_{goal}(\bm{\tau})+J_{smooth}(\bm{\tau})$ (1)
@@
$f_{contact}(\mathbf{s}_{t})=0,\quad f_{kinematics}(\mathbf{s}_{t},\mathbf{s}_{t+1})=0$
@@
$f_{friction}(\mathbf{s}_{t},\mathbf{u}_{t})\leq 0,\quad f_{min\_f}(\bm{u}_{t})\leq 0$
where the state space is defined as $\mathbf{s}_{t}:=\{\mathbf{q}_{t},\mathbf{o}_{t}\}$. The action space is defined as $\mathbf{u}_{t}:=\{\hat{\mathbf{u}}_{t},\mathbf{f}_{1,t},\cdots,\mathbf{f}_{N,t},\mathbf{f}_{e,t}\}$, where $\hat{\mathbf{u}}_{t}$ is the delta action, $\mathbf{f}_{i,t}$ and $\mathbf{f}_{e,t}$ are the forces that the $i$th finger and the environment apply to the object respectively. The objective $J_{goal}$ incentivizes the robot to manipulate the object to the desired pose, and $J_{smooth}$ is defined as $J_{smooth}=\sum_{t=1}^{T}||\mathbf{s}_{t}-\mathbf{s}_{t-1}||_{2}^{2}$. $f_{contact}$ ensures the fingers are always in contact with the object. $f_{kinematics}$, $f_{balance}$, and $f_{friction}$ are kinematics constraints, wrench balance constraints, and friction constraints, which are explained in the section below. $f_{min\_f}$ represents the constraint that ensures the force norm remains above a specified threshold, addressing the need to overcome joint friction and maintain contact in hardware experiment [35]. Our primary contribution lies in the formulation of $f_{contact}$ and $f_{kinematics}$, utilizing SDF and a sampling-based method for estimating constraints and their gradients. We use CSVTO [9] to solve the optimization.
@@ IV-A1 Contact Constraints
The constraints are formulated as $\Phi_{i}(\mathbf{q}_{t},\mathbf{o}_{t})=0$, where $\Phi_{i}$ returns the distance between the $i$th finger and the object. The distance between two shapes is defined as the minimal distance between any two points sampled from their surfaces. As robot fingers can have non-primitive geometries, estimating the constraints requires parameterizing the geometries. Additionally, the estimation must be differentiable to allow the gradients of the constraints to be used for updating the trajectory. We will further discuss the geometry processing in Sec. IV-B.
@@ IV-A2 Kinematics Constraint
$J_{\mathbf{c}_{i,t}}(\mathbf{q}_{t})\dot{\mathbf{q}}_{t}=\bm{\omega}_{t}\times\mathbf{r}(\mathbf{c}_{i,t},\mathbf{o}_{t})+\dot{\mathbf{x}}_{t},\vspace{-0.1cm}$ (2)
@@
However, having both the kinematics constraint and the contact constraint may over-constrain the problem: approximating velocities by finite differencing positions, both the contact constraint (1D constraint) and the kinematics constraint (3D constraint) specify the contact point location (3D variable), adding up to a 4D constraint. Thus, we project the 3D kinematics constraints into the 2D tangent space of the contacts: $R(\mathbf{n}_{i,t})(J_{\mathbf{c}_{i,t}}(\mathbf{q}_{t})\dot{\mathbf{q}}_{t}-\bm{\omega}_{t}\times\mathbf{r}(\mathbf{c}_{i,t},\mathbf{o}_{t})-\dot{\mathbf{x}}_{t}),$ where $R(\mathbf{n}_{i,t})$ is the projection matrix based on the contact normal $\mathbf{n}_{i,t}$ of the $i$th finger. See Fig. 2b for visualization.
Similar to contact constraints, estimating $\mathbf{c}_{i,t}$, $J_{\mathbf{c}_{i,t}}$, $\mathbf{n}_{i,t}$ and their gradients also requires parameterizing the geometries, which is discussed in Sec. IV-B.